# The Codecfake Dataset and Countermeasures for the Universally Detection of Deepfake Audio

### v2

@@ VI-C Impact of Codec Settings
To further validate different codec parameters setting will have influence performance of ADD model, we tested various configurations of C3 and C4 using a codec-trained ADD model as shown in Table VI. It can be observed that compared to the original settings of C3 and C4, using different parameters to generate codec-based audio does not significantly impact the performance of the anti-spoofing system. Especially for C3-FuncCodec, the EERs of C3-1 to C3-4 across the three ADD systems are almost identical. For C4-Encodec, it can be observed that from C4-1 to C4-3, the EER of the ADD model gradually increases. This indicates that using higher bps and quantizers in codec synthesis system tends to produce more authentic audio, making it more challenging for the ADD model to distinguish.
TABLE VIII: EER (%) results for ADD methods co-trained by 19LA training set and Codecfake training set.
Model 19LA IWA C1 C2 C3 C4 C5 C6 C7 CAVG AVG
@@
VI-D Test on ALM-based audio
In practice, ALM-based audio encompasses not only the reconstruction by neural codec method but also manipulation in content. To assess whether ADD models trained on the codecfake dataset can effectively detect ALM-based audio, we conducted experiments as shown in the Table VII. At first, we observed that employing solely vocoder-trained ADD models yielded satisfactory outcomes for ALM-based audio, particularly evident in evaluations on A1 and A2, where models such as W2V2-AASIST achieved EERs below 1%. However, substantial performance deterioration was evident in evaluations on the condition A3, attributable to OOD issues. The ALM test condition A3 is characterized by sounds with event descriptions that deviate from the distribution of speech. After using Codecfake to train ADD model, performance across A1-A3 are improved, with W2V2-AASIST achieving the lowest average EER of 11.388%, representing a decrease of 4.189% compared to the vocoder-trained ADD models. However, it is noteworthy that while the EER for A3 decreased compared to the vocoder-trained ADD models, it still remains relatively high.
@@ VI-D Test on ALM-based audio
VI-E Countermeasure for generalized ADD method
@@
Through the experiments described above, we observed a phenomenon: while training in different domains yields optimal performance in their respective test domains, it leads to performance degradation in cross-domain test scenarios. In this section, we propose a co-training approach using vocoder-based audio and codec-based audio to ensure that the detection model performs well across all scenarios. Furthermore, during the co-training process, we employed the proposed CSAM method. This approach optimizes the ERM while learning flat minima, avoiding the domain ascent bias problem that may arise during the co-train situation in SAM. Experiment result is shown in Table VIII, where SAM is original implementation, ASAM[75] is a modified SAM approach to improve generalization performance by parameter re-scaling, CSAM is our proposed approach. Firstly, the performance of the co-trained ADD model has improved compared to both vocoder-trained ADD model and codec-trained ADD model under the testing conditions. For instance, the top-performing model, W2V2-AASIST, achieved a CAVG of 0.122% and an AVG EER of 0.963%, compared to the vocoder-trained W2V2-AASIST with a CAVG of 41.583% and an AVG EER of 34.991%, and the codec-trained W2V2-AASIST with a CAVG of 0.177% and an AVG EER of 1.628%, indicating a decrease in EER. Then, to further enhance the model’s generalizability, we trained all model in the co-training scenario with SAM, ASAM, and CSAM strategy. After employing the original SAM, it was observed that apart from an increase in EER on the IWA dataset and the C7 unseen test set, there was a decrease in EER across all other test scenarios. Particularly noteworthy were the remarkable scores of 0% EER obtained under the C3 and C4 test conditions. After further employing CSAM, compared to original SAM, although there was a slight increase in EER on the visible test condition, there was a further improvement in the performance on unseen OOD test conditions. Specifically, using W2V2-AASIST (CSAM) resulted in a decrease in EER from 7.483% to 4.689% on the IWA test condition and from 0.627% to 0.431% on the unseen C7 test set, compared to W2V2-AASIST (SAM). Finally, W2V2-AASIST (CSAM) achieved the lowest average EER, with a CAVG of 0.077% and an AVG of 0.616%.
@@
To further investigate whether SAM benefits the model training, we plotted the loss curves of co-trained W2V2-AASIST during training and validation as shown in the Figure 6. From Figure 6(a), we observed that the convergence process of loss is similar across different SAM methods. However, CSAM exhibits lower loss values as early as the first epoch, indicating that the CSAM approach can accelerate the convergence speed of ERM. For the validation process, whether SAM, ASAM, or CSAM is applied, there is a noticeable decrease in loss compared to when SAM is not used, especially in the first epoch. Furthermore, for epochs 5-10, we observe that without SAM, the seventh epoch is optimal for the model, while the SAM series consistently optimizes the validation loss, achieving the best model in the ninth epoch. When training a ADD model, even if the 19LA task runs for 100 epochs, the optimal checkpoint often emerges around epochs 20-30. Subsequent reductions in the learning rate are insufficient to learn a more generalized model on the validation set. However, SAM optimizes ERM while simultaneously learning flat minima, allowing each epoch to continuously seek the most generalizable solution. In the validation set, CSAM achieves the lowest loss in the ninth epoch, corresponding to its best performance on the test set.
@@ VII Limitation and Future direction
Generalized ADD methods. The current ADD methods, such as W2V2-AASIST, exhibit a significant performance decline in OOD scenarios. In this paper, we attempted to achieve a generalized ADD model using co-training and CSAM approaches. However, there remains significant for improvement in OOD scenarios such as IWA and A3. Designing a generalized approach that can detect both vocoder-based audio and ALM-based audio is currently a crucial challenge.

### v3

@@ VI-C Impact of Codec Settings
To further validate different codec parameters setting will have influence performance of ADD model, we tested various configurations of C3 and C4 using a codec-trained ADD model as shown in Table VII. It can be observed that compared to the original settings of C3 and C4, using different parameters to generate codec-based audio does not significantly impact the performance of the anti-spoofing system. Especially for C3-FuncCodec, the EERs of C3-1 to C3-4 across the three ADD systems are almost identical. For C4-Encodec, it can be observed that from C4-1 to C4-3, the EER of the ADD model gradually increases. This indicates that using higher bps and quantizers in codec synthesis system tends to produce more authentic audio, making it more challenging for the ADD model to distinguish.
VI-D Impact of language setting
Codecfake is generated based on two source languages: Chinese from AISHELL and English from VCTK. In this section, we explore the impact of language on ADD models. We conducted the following experiments, first dividing the Codecfake training set, development set, and evaluation set based on AISHELL and VCTK. The experiments are categorized into intra-language testing and cross-language testing. In intra-language testing, both the training and testing languages are kept consistent. While in cross-language testing, the training and testing languages differ, with the experiment results presented in Table VIII.
From the experiments results, we can draw several conclusions. First, cross-language testing results in a significant performance drop compared to intra-language testing. For example, the EER for English training with Chinese testing worsens by 3.669% compared to English training with English testing, and the EER for Chinese training with English testing worsens by 1.536% compared to Chinese training with Chinese testing. Second, the ADD model trained solely on English outperforms the model trained on Chinese. When testing on the complete set of Codecfake test conditions, the model trained in English achieves an EER of 2.770%, while the model trained in Chinese achieves an EER of 7.024%. Additionally, in intra-language testing, training and testing in English achieve an EER of 0.205%, which is the best result in language testing experiments. Finally, the performance of multilingual training surpasses that of training solely on one language. By utilizing the entire multilingual Codecfake training set, the CAVG achieves an EER of 0.177%, significantly lower than when using only a single language for training. This demonstrates the advantages of the multilingual design of Codecfake.
TABLE X: EER (%) results for ADD methods co-trained by 19LA training set and Codecfake training set.
Model 19LA ITW C1 C2 C3 C4 C5 C6 C7 CAVG AVG
@@ VI-D Impact of language setting
VI-E Test on ALM-based audio
In practice, ALM-based audio encompasses not only the reconstruction by neural codec method but also manipulation in content. To assess whether ADD models trained on the codecfake dataset can effectively detect ALM-based audio, we conducted experiments as shown in the Table IX. At first, we observed that employing solely vocoder-trained ADD models yielded satisfactory outcomes for ALM-based audio, particularly evident in evaluations on A1 and A2, where models such as W2V2-AASIST achieved EERs below 1%. However, substantial performance deterioration was evident in evaluations on the condition A3, attributable to OOD issues. The ALM test condition A3 is characterized by sounds with event descriptions that deviate from the distribution of speech. After using Codecfake to train ADD model, performance across A1-A3 are improved, with W2V2-AASIST achieving the lowest average EER of 11.388%, representing a decrease of 4.189% compared to the vocoder-trained ADD models. However, it is noteworthy that while the EER for A3 decreased compared to the vocoder-trained ADD models, it still remains relatively high.
@@ VI-E Test on ALM-based audio
VI-F Countermeasure for generalized ADD method
@@
Through the experiments described above, we observed a phenomenon: while training in different domains yields optimal performance in their respective test domains, it leads to performance degradation in cross-domain test scenarios. In this section, we propose a co-training approach using vocoder-based audio and codec-based audio to ensure that the detection model performs well across all scenarios. Furthermore, during the co-training process, we employed the proposed CSAM method. This approach optimizes the ERM while learning flat minima, avoiding the domain ascent bias problem that may arise during the co-train situation in SAM. Experiment result is shown in Table X, where SAM is original implementation, ASAM[83] is a modified SAM approach to improve generalization performance by parameter re-scaling, CSAM is our proposed approach. Firstly, the performance of the co-trained ADD model has improved compared to both vocoder-trained ADD model and codec-trained ADD model under the testing conditions. For instance, the top-performing model, W2V2-AASIST, achieved a CAVG of 0.122% and an AVG EER of 0.963%, compared to the vocoder-trained W2V2-AASIST with a CAVG of 41.583% and an AVG EER of 34.991%, and the codec-trained W2V2-AASIST with a CAVG of 0.177% and an AVG EER of 1.628%, indicating a decrease in EER. Then, to further enhance the model’s generalizability, we trained all model in the co-training scenario with SAM, ASAM, and CSAM strategy. After employing the original SAM, it was observed that apart from an increase in EER on the ITW dataset and the C7 unseen test set, there was a decrease in EER across all other test scenarios. Particularly noteworthy were the remarkable scores of 0% EER obtained under the C3 and C4 test conditions. After further employing CSAM, compared to original SAM, although there was a slight increase in EER on the visible test condition, there was a further improvement in the performance on unseen OOD test conditions. This is because SAM tends to overfit on a specific ID domain during multi-domain training, whereas CSAM computes each mini-batch across all domains, which can significantly enhances OOD performance. Specifically, using W2V2-AASIST (CSAM) resulted in a decrease in EER from 7.483% to 4.689% on the ITW test condition and from 0.627% to 0.431% on the unseen C7 test set, compared to W2V2-AASIST (SAM). Finally, W2V2-AASIST (CSAM) achieved the lowest average EER, with a CAVG of 0.077% and an AVG of 0.616%.
@@
To further investigate whether SAM benefits the model training, we plotted the loss curves of co-trained W2V2-AASIST during training and validation as shown in the Figure 6. From Figure 6(a), we observed that the convergence process of loss is similar across different SAM methods. However, CSAM exhibits lower loss values as early as the first epoch, indicating that the CSAM approach can accelerate the convergence speed of ERM. From Figure 6(b), in the validation process, whether SAM, ASAM, or CSAM is applied, there is a noticeable decrease in loss compared to when SAM is not used, especially in the first epoch. Furthermore, for epochs 5-10, we observe that without SAM, the seventh epoch is optimal for the model, while the SAM series consistently optimizes the validation loss, achieving the best model in the ninth epoch. When training a ADD model, even if the 19LA task runs for 100 epochs, the optimal checkpoint often emerges around epochs 20-30. Subsequent reductions in the learning rate are insufficient to learn a more generalized model on the validation set. However, SAM optimizes ERM while simultaneously learning flat minima, allowing each epoch to continuously seek the most generalizable solution. In the validation set, CSAM achieves the lowest loss in the ninth epoch, corresponding to its best performance on the test set.
@@ VII Limitation and Future direction
Generalized ADD methods. The current ADD methods, such as W2V2-AASIST, exhibit a significant performance decline in OOD scenarios. In this paper, we attempted to achieve a generalized ADD model using co-training and CSAM approaches. However, there remains significant for improvement in OOD scenarios such as ITW and A3. Designing a generalized approach that can detect both vocoder-based audio and ALM-based audio is currently a crucial challenge.